# A Comprehensive Taxonomy of Negation for NLP and Neural Retrievers

### v1

@@ 6.4.1 Evaluation on NevIR
As shown in Table 1 and in Figure 13 in Appendix A.8.1, fine-tuning ColBERT and MultiQA on our synthetic dataset yields an immediate performance gain on the NevIR development set, however peaking while fine-tuning on NevIR train reaches higher performance in the last epoch. This is to be expected as for the synthetic data we evaluate OOD. To assess in-distribution performance, we apply mixed fine-tuning by combining the two datasets and shuffling the data. The model achieves high performance significantly faster than when simply fine-tuned on NevIR, giving the overall best performance. Mistral shows the same behaviour with mixed fine-tuning. This supports hypothesis H4, that our synthetically generated dataset helps in capturing negation.
@@ Appendix A Appendix
This appendix offers further material that supports the study. It is organised as follows: Appendix A.1 defines the properties of negation that are briefly referenced in the study. Appendix A.2 gives an example in an information retrieval style for each type of negation present in the taxonomy, alongside further definitions of exceptors and typed lambda calculus. Appendix A.3 lists all the prompts used to generate the datasets. Appendix A.4 mentions use cases that we do not explicitly account for in this study, although they are interesting to study. A.5 lists details into applying the categorization mechanism on the ExcluIR dataset. Appendix A.6 includes the survey that the human annotators completed to perform a qualitative evaluation of the generated data. Appendix A.7 contains a table with specifications of all models used in this study: architecture, training objective, training dataset, size and tokenizer - these specifications influence the model’s performance on encoding negation. Appendix A.8 contains the results of evaluating the models against the controlled generated dataset and the ExcluIR data. Finally, Appendix A.6.1 offers a statistical analysis of the annotator’s answers.
@@ A.6.1 Statistical analysis on annotation results
A.7 Models
| BM25 | Sparse | Retrieval | N/A | N/A | N/A |
| DPR [29] | Bi-Encoder | Retrieval | NQ | 219M | BERT |
| coCondenser [14] | Bi-Encoder | Retrieval | MSMarco | 110M | BERT |
| Dragon [36] | Bi-Encoder | Retrieval | MS MARCO | N/A | BERT |
| Sentence-T5 | Dual Encoder | Sentence Similarity | NLI | 220M | T5 |
| ColBERTv1 [32] | Late Interaction | Retrieval | MSMarco | 110M | BERT |
| ColBERTv2 [58] | Late Interaction | Retrieval | MSMarco | 110M | BERT |
| MonoT5 Base [51] | Crossencoder | Ranking | MSMarco | 223M | T5 |
| MonoT5 Large [51] | Crossencoder | Ranking | MSMarco | 737M | T5 |
| MonoT5 3B [51] | Crossencoder | Ranking | MSMarco | 2.85B | T5 |
| qnli-electra-base | Crossencoder | NLI | QNLI | 110M | ELECTRA |
| Mistral-7B-Instruct [26] | Transformer | NTP | Crawled | 7B | BPE |
| Llama-3.1-3B-Instruct [16] | Transformer | NTP | Crawled | 7B | Llama |
| Llama-3.2-8B-Instruct [16] | Transformer | NTP | Crawled | 7B | Llama |
A.8 Results

### v2

@@ 6.4.1 Evaluation on NevIR
As shown in Table 2 and in Figure 13 in Appendix A.7.1, fine-tuning ColBERT and MultiQA on our synthetic dataset yields an immediate performance gain on the NevIR development set, however peaking while fine-tuning on NevIR train reaches higher performance in the last epoch. This is to be expected as for the synthetic data we evaluate OOD. To assess in-distribution performance, we apply mixed fine-tuning by combining the two datasets and shuffling the data. The model achieves high performance significantly faster than when simply fine-tuned on NevIR, giving the overall best performance. Mistral shows the same behaviour with mixed fine-tuning. This supports hypothesis H4, that our synthetically generated dataset helps in capturing negation. Overall, we notice that fine-tuning on our synthetic data brings a quick performance boost against the NevIR dev and test sets, indicating that our proposed datasets capture the notion of negation.
@@ 7 Conclusion
Acknowledgments
The evaluation of our generated data was done through LabelStudio. Moreover, we acknowledge our colleagues who helped with human evaluation and annotation: Panagiotis Eustratiadis, Jasmin Kareem, Clara Rus, David Vos, Maria Heuss, Lu Zhang, and Catherine Chen. We also want to acknowledge Maria Aloni, who offered help and feedback for our linguistic study.
This research was (partially) supported by the Dutch Research Council (NWO), under project numbers 024.004.022, NWA.1389.20.183, and KICH3.LTP.20.006, the European Union under grant agreements No. 101070212 (FINDHR) and No. 101201510 (UNITE), and Ahold Delhaize. Views and opinions expressed are those of the author(s) only and do not necessarily reflect those of their respective employers, funders and/or granting authorities.
@@ Appendix A Appendix
This appendix offers further material that supports the study. It is organised as follows: Appendix A.1 defines the properties of negation that are briefly referenced in the study. Appendix A.2 gives an example in an information retrieval style for each type of negation present in the taxonomy, alongside further definitions of exceptors and typed lambda calculus. Appendix A.3 lists all the prompts used to generate the datasets. Appendix A.4 mentions use cases that we do not explicitly account for in this study, although they are interesting to study. A.5 lists details into applying the categorization mechanism on the ExcluIR dataset. Appendix A.6 includes the survey that the human annotators completed to perform a qualitative evaluation of the generated data. Appendix A.7 contains the results of evaluating the models against the controlled generated dataset and the ExcluIR data. Finally, Appendix A.6.1 offers a statistical analysis of the annotator’s answers.
@@ A.6.1 Statistical analysis on annotation results
A.7 Results